# Special Weingarten surfaces with planar convex boundary

### v1

@@ Introduction
We point out that, in most of the cited works, $W$ is expressed in terms of the mean curvature $H$ and the Gaussian curvature $K$, resulting in the symmetry of $W$ with respect to the principal curvatures. We do not make this assumption and the absence of symmetry leads to the non-differentiability of the underlying partial differential equation at the umbilical points of the SW-surface. Then, as we need a maximum principle, we will rely on a result by Alexandrov [Al], where the symmetry of the elliptic operator is not assumed. We will discuss this issue in Section 1.3.
@@
We are going to prove that, under suitable conditions, such surface is topologically a closed disk. Moreover it is the union of vertical and cylindrical graphs. This is a generalization of the analogous result by A. Ros & H. Rosenberg for constant mean curvature surfaces in ${\mathbb{R}}^{3}$ [RoRo, Theorem 2]. However, our proof follows the strategy of [Se, NePi, NePiRu], which do not make use of techniques as curvature estimates and compactness theorems, which are not known in the context of SW-surfaces.
@@
The first step towards a generalization of Ros-Rosenberg Theorem for SW-surfaces is to notice that for any set $\mathcal{W}_{g}$ of either minimal-type or CGC-type, without any further restriction on the function $g$, the proof is a straightforward consequence of the Convex Hull Property and the Gauss-Bonnet Theorem, respectively (see Remark 5.2). On the other hand, when $\mathcal{W}_{g}$ consists of SW-surfaces of CMC-type, the situation is not elementary, and we present two setting where applying our approach.
First, we consider those sets $\mathcal{W}_{g}$ of CMC-type whose corresponding surfaces are related by some homothety of $\mathbb{R}^{3}$ and we define an equivalence relation among the functions $g$. Then if a function $g$ satisfies an additional but natural global analytic assumption, we get Ros-Rosenberg result for every surfaces obtained from surfaces in $\mathcal{W}_{g}$ after a suitable homothety (see Theorem 5.1). This first case includes, in particular, any set $\mathcal{W}_{g}$ of uniformly elliptic Weingarten surfaces, that is, when there exist constants $c_{0}<C_{0}<0$ such that $c_{0}\leq g^{\prime}(t)\leq C_{0}$ for all $t\in I_{g}=[\alpha,+\infty)$.
In the latter setting, we do not place any global assumption on the behavior of the function $g$, but we impose instead a mild geometric condition involving the curvatures of the sphere and the cylinder belonging to $\mathcal{W}_{g}$ (see Theorem 5.6). As a consequence, we enlarge the possible sets $\mathcal{W}_{g}$ of CMC-type for which Ros-Rosenberg result holds. For example, we can consider a function $g$ whose whose graph is asymptotic to a vertical line, that clearly does not satisfy the assumptions of Theorem 5.1. It is also important to notice that Theorem 5.1 and Thorem 5.6 are not consequence one of the other, but they are complementary, although their proofs shares most of the arguments.
The paper is organized as follows.
Section 1 is devoted to preliminaries. In particular, we give the precise definition of SW-surface and we introduce the curvature diagram of a surface. Moreover, we define an equivalence relation for sets $\mathcal{W}_{g}$ of CMC-type related by a homothety of ${\mathbb{R}}^{3}$ and we recall the tangency principle for SW-surfaces. In Section 2, we present a class of Special Weingarten catenoids, governed by a linear relation of the form $\kappa_{2}=m_{0}\kappa_{1}$, that we intend to use as barriers in the proofs of theorems 5.1 and 5.6. In Section 3, we describe the geometric properties of a classic curve called limaçon of Pascal, that will be useful in our proofs. In Section 4, we prove an important auxiliary lemma. Finally, in Section 5, we present our two generalizations of Ros-Rosenberg Theorem for SW-surfaces (Theorem 5.1 and Theorem 5.6) as well as their proofs.
@@ 1.1 The curvature diagram and the definition of SW-surfaces
Notice that for a set of SW-surfaces given in terms of Definition 0.1 satisfying condition (0.2), it is possible that the set $W^{-1}(\{0\})$ may contain more than one connected component. In general, surfaces belonging to different connected component are topologically and geometrically different and they are not comparable by the maximum principle, hence they are not suitable to be considered together. As an example, consider the simple case where $W(\kappa_{1},\kappa_{2})=\kappa_{1}+\kappa_{2}-\kappa_{1}\kappa_{2}$. Then $W^{-1}(\{0\})$ has two connected components which are contained in different branches of a hyperbola (see Fig. 2, left). When $\kappa_{1}>1$, the branch of this hyperbola includes round spheres of radius $1/2$, while for $\kappa_{1}<1$ the branch of this hyperbola includes planes.
@@
A Special Weingarten surface (SW-surface, for short) is an oriented surface $\Sigma$ immersed in $\mathbb{R}^{3}$ whose principal curvatures $\kappa_{1}\geq\kappa_{2}$ satisfy at every point of $\Sigma$ a relation of the type $\kappa_{2}=g(\kappa_{1})$, for some $g\in C^{1}(I_{g})$ with $g^{\prime}(t)<0$ everywhere, where $I_{g}=[\alpha,b)$ for some $b\in\mathbb{R}\cup\{+\infty\}$ and $g(\alpha)=\alpha\neq 0$.
@@ 1.2 An equivalence relation among Special Weingarten surfaces
1.3 The maximum principle for Special Weingarten surfaces
@@ 2 SW-catenoids
Special Weingarten surfaces with rotational symmetry are described in details and classified in [FeMi]. The minimal-type case was previously treated in [SaTo1] and more recently in [LoPa]. Recall that, if $\Sigma$ is a surface in $\mathcal{W}_{g}$ for a given $g$ with $\alpha=0$, then $\Sigma$ is called of minimal-type. Regular surfaces invariant by rotations in a set of SW-surfaces $\mathcal{W}_{g}$ of minimal-type are called special catenoids or SW-catenoids.
@@
Let us consider an SW-catenoid $G\in\mathcal{W}_{g}$ where $g(t)=m_{0}t$, $m_{0}<0$, which is inward oriented. In this case, we call $G$ a linear catenoid. After an isometry of $\mathbb{R}^{3}$ we can assume that the revolution axis is the $z$-axis, and the neck is the horizontal circle $C_{0}$ of radius $r_{0}>0$ centered at the origin. The part of $G$ above the plane $\{z=0\}$ can be parametrized as
@@ 3 The limaçon of Pascal
Denoting by $C$ and $c$ respectively the center and the radius of $\mathcal{C}$, it is easy to prove that, up to isometries of the Euclidean plane, $\mathcal{L}$ is uniquely determined by two positive parameters: $a=\mbox{dist}(A,C)$ and $c$. Moreover $\mathcal{L}$ depends continuously by $a$ and $c$.
@@ 4 The graph Lemma
We denote by $D(p,r)$ the horizontal closed disk of radius $r$ and center $p.$ Morover, for any interval $I\subset\mathbb{R}$, we denote by $D(p,r)\times I$ the intersection of the solid cylinder on $D(p,r)$ with the closed slab $\{(x,y,z)\in\mathbb{R}^{3};z\in I\}$.
Let us prove a result that will be crucial in what follows.
@@
If $\Omega$ is the planar domain bounded by $\Gamma,$ then there exists $r=r(\Gamma)$ and $p\in\Omega$ such that $\Sigma\cap(D(p,r)\times[0,+\infty))$ is a vertical graph over $D(p,r)$, with
@@
We denote by $\mathcal{O}$ the open region of $\mathbb{R}^{3}$ bounded by $\Sigma\cup\Omega$.
@@
In order to prove the statement, we apply Alexandrov reflection method with horizontal hyperplanes $P(t):=\{z=t\}$ coming down from above. Denote by $h_{\Sigma}$ the height of $\Sigma$ above $P(0)$. The plane $P(t)$, $t>h_{\Sigma}$, does not intersect $\Sigma$. Then we let $t$ decrease. When $t<h_{\Sigma}$, we reflect the part of $\Sigma$ above $P(t)$ until there is a first contact point between $\Sigma$ and its reflection. If such a contact point does not occur for any $t>0$, then $\Sigma$ is a vertical graph over $\Omega$ and we can choose $r(\Gamma)=1/\Lambda$ and $p=q+(1/\Lambda)n(q)$, where $q$ is a point of $\Gamma$ and $n(q)$ is the inward pointing unit normal vector of $\Gamma$ at $q$.
@@
Otherwise, there will be a $20<t_{0}<h_{\Sigma}/$ 2 such that the reflected surface touches $\Sigma$ for the first time. If the intersection occurs at an interior point of $\Sigma$ we would have a contradiction with the maximum principle, hence a first touching point must belong to $\Gamma$. Let $q$ be one of such points. Then the line $\{q\}\times(0,+\infty)$ intersects $\Sigma$ exactly once, and $\{q\}\times(0,2t_{0})$ is contained in $\mathcal{O}$, as $t_{0}<h_{M}/2$. Note that the portion of $\Sigma$ above $P(t_{0})$ is a graph.
@@
We now perform Alexandrov’s reflections with respect to vertical planes. Let $Q$ be one of such planes. Since $\Sigma$ is compact, we can assume that $Q\cap\Sigma=\emptyset$. Translate horizontally $Q$ towards $\Sigma$. By abuse of notation, we call again $Q$ any parallel translation of the initial plane. When $Q$ touches $\Sigma$ for the first time, keep moving $Q$ and start reflecting through $Q$ the part of $\Sigma$ left behind $Q$. In order not to have a contradiction with the maximum principle, we can continue this procedure with no contact points between $\Sigma$ and its reflection until $Q$ enters $\Gamma$ at distance at least $1/\Lambda$ from it.
We can avoid the dependence on the contact point $q$ by stopping reflecting when $Q$ is tangent to $\mathcal{C}$, where $\mathcal{C}$ is defined as follows. Denote by $\mathcal{C}^{\prime}$ the horizontal circle of radius $1/\lambda$, tangent to $\Gamma$ at $q$, and enclosing $\Gamma$. Then $\mathcal{C}$ is the circle with the same center as that of $\mathcal{C}^{\prime}$ and radius equal to $1/\lambda-1/\Lambda$.
@@
The set of the reflections of $q$ through any vertical plane tangent to $\mathcal{C}$ is a limaçon of Pascal $\mathcal{L}$ in the plane $P$, with respect to the circle $\mathcal{C}$ and base point $q$. The parameters of $\mathcal{L}$ are $a=1/\lambda$ and $c=1/\lambda-1/\Lambda$. As $a>c$, then $\mathcal{L}$ has two loops. By item (2) of Lemma 3.3, the smaller loop of $\mathcal{L}$ is contained in the circle of radius $a-c=1/\Lambda$ tangent to $\Gamma$ at $q$ hence it is contained in $\Omega\cup\{q\}$.
Furthermore, since $\{q\}\times(0,+\infty)$ intersects $\Sigma$ in exactly one point, then the same holds true for any point in the closed domain bounded by the smaller loop of $\mathcal{L}$. Indeed, we know that the part of $\Sigma$ above the plane $P(t_{0})$ is a vertical graph, so we only need to prove that for any point $\widetilde{q}$ in the smaller loop of $\mathcal{L}$, the rectangle =$\mathcal{R}=\{(1-s)q+s\widetilde{q},s\in[0,1]\}\times(0,t_{0})$ lies inside $\mathcal{O}$. For a such point $\widetilde{q}$ there exists a vertical plane $\widetilde{Q}$ tangent to $\mathcal{C}\times\mathbb{R}$ for which $\widetilde{q}$ is the reflection of $q$ across $\widetilde{Q}$. Let $\widetilde{Q}(s)$ denote the translation of $\widetilde{Q}$ by the vector $s(q-\widetilde{q})$. Notice that by construction of $\mathcal{C}$, for any $s\geq 0$ the plane $\widetilde{Q}(s)$ enters $\Gamma$ at most $1/\Lambda$. Then, by the Alexandrov reflection method, for any $s\geq 0$ the part of $\Sigma$ in the same side of $q$ relative to $\widetilde{Q}(s)$ is a horizontal graph and its reflection across $\widetilde{Q}(s)$ lies in $\mathcal{O}$. In particular, since the line segment $\{q\}\times(0,t_{0})$ is contained in $\mathcal{O}$, its reflection across $\widetilde{Q}(s)$ also lies inside $\mathcal{O}$, for any $0\leq s\leq 1$. This proves that the rectangle $\mathcal{R}$ lies inside $\mathcal{O}$.
@@
Finally, the conclusion of the statement follows from the fact that the closed domain bounded by the smaller loop of $\mathcal{L}$ contains a closed disk $D(p,r)$, for some $p\in\Omega$ and $r=r(\Gamma)$ greater or equal than $\sqrt{\frac{(a-c)^{3}}{a}}=\sqrt{\frac{\lambda}{\Lambda^{3}}}$, by item (1) of Lemma 3.3.

### v2

@@ Introduction
We point out that, in most of the cited works, $W$ is expressed in terms of the mean curvature $H$ and the Gaussian curvature $K$, resulting in the symmetry of $W$ with respect to the principal curvatures. We do not make this assumption and the absence of symmetry leads to the non-differentiability of the underlying partial differential equation at the umbilical points of the SW-surface. Then, as we need a tangency principle, we will rely on a result by Alexandrov [Al], where the symmetry of the elliptic operator is not assumed. We will discuss this issue in Section 1.3.
@@
We are going to prove that, under suitable conditions, such surface is topologically a closed disk. Moreover it is the union of vertical and possibly cylindrical graphs. This is a generalization of the following analogous result for constant mean curvature surfaces in ${\mathbb{R}}^{3}$ proved by A. Ros & H. Rosenberg.
Theorem.
[RoRo, Theorem $2$] Let $\Gamma\subset P$ be a strictly convex curve. There is an $H(\Gamma)>0$, depending only on the extreme values of the curvature of $\Gamma$, such that whenever $M\subset\mathbb{R}^{3}_{+}$ is asurface with constant mean curvature $H$ bounded by $\Gamma$, with $0<H<H(\Gamma)$, then $M$ is topologically a disk and either $M$ is a graph over the domain $\Omega$ bounded by $\Gamma$ or $N=M\cap(\Omega\times]0,\infty[)$ is a graph over $\Omega$ and $M-N$ is a graph over a subannulus of $\Gamma\times]0,\infty[$, with respect to the lines normal to $\Gamma\times]0,\infty[$.
Notice that our proof follows the strategy of [Se, NePi, NePiRu], which do not make use of techniques as curvature estimates and compactness theorems, which are not known in the context of SW-surfaces.
@@
The first step towards a generalization of Ros-Rosenberg Theorem for SW-surfaces is to notice that for any set $\mathcal{W}_{g}$ of either minimal-type or CGC-type, without any further restriction on the function $g$, the proof is a straightforward consequence of the Convex Hull Property and the Gauss-Bonnet Theorem, respectively (see Remark 5.2). On the other hand, when $\mathcal{W}_{g}$ consists of SW-surfaces of CMC-type, the situation is not elementary, and we present two settings where applying our approach.
First, we consider those sets $\mathcal{W}_{g}$ of CMC-type whose corresponding surfaces are related by some homothety of $\mathbb{R}^{3}$ and we define an equivalence relation among the functions $g$. If a function $g$ satisfies an additional but natural global analytic assumption, we get Ros-Rosenberg result for all surfaces obtained from surfaces in $\mathcal{W}_{g}$ after a suitable homothety (see Theorem 5.1). This first case includes, in particular, any set $\mathcal{W}_{g}$ of uniformly elliptic Weingarten surfaces, that is, when there exist constants $c_{0}<C_{0}<0$ such that $c_{0}\leq g^{\prime}(t)\leq C_{0}$ for all $t\in I_{g}=[\alpha,+\infty)$.
In the latter setting, we do not place any global assumption on the behavior of the function $g$, but we impose instead a mild geometric condition involving the curvatures of the sphere and the cylinder belonging to $\mathcal{W}_{g}$ (see Theorem 5.6). As a consequence, we enlarge the possible sets $\mathcal{W}_{g}$ of CMC-type for which Ros-Rosenberg result holds. For example, we can consider a function $g$ whose whose graph is asymptotic to a vertical line, that clearly does not satisfy the assumptions of Theorem 5.1. It is also important to notice that Theorem 5.1 and Theorem 5.6 are not consequence one of the other, but they are complementary, although their proofs shares most of the arguments.
The paper is organized as follows. Section 1 is devoted to preliminaries. In particular, we give the precise definition of SW-surface and we introduce the curvature diagram of a surface. Moreover, we define an equivalence relation for sets $\mathcal{W}_{g}$ of CMC-type related by a homothety of ${\mathbb{R}}^{3}$ and we recall the tangency principle for SW-surfaces. In Section 2, we present a class of Special Weingarten catenoids, governed by a linear relation of the form $\kappa_{2}=m_{0}\kappa_{1}$, that we intend to use as barriers in the proofs of theorems 5.1 and 5.6. In Section 3, we describe the geometric properties of a classic curve called limaçon of Pascal, that will be useful in our proofs. In Section 4, we prove an important auxiliary lemma. Finally, in Section 5, we present our two generalizations of Ros-Rosenberg Theorem for SW-surfaces (Theorem 5.1 and Theorem 5.6) as well as their proofs.
@@ 1.1 The curvature diagram and the definition of SW-surfaces
Notice that for a set of SW-surfaces given in terms of Definition 0.1 satisfying condition (0.2), it is possible that the set $W^{-1}(\{0\})$ may contain more than one connected component. In general, surfaces belonging to different connected component are topologically and geometrically different and they are not comparable by the tangency principle, hence they are not suitable to be considered together. As an example, consider the simple case where $W(\kappa_{1},\kappa_{2})=\kappa_{1}+\kappa_{2}-\kappa_{1}\kappa_{2}$. Then $W^{-1}(\{0\})$ has two connected components which are contained in different branches of a hyperbola (see Fig. 2, left). When $\kappa_{1}>1$, the branch of this hyperbola includes round spheres of radius $1/2$, while for $\kappa_{1}<1$ the branch of this hyperbola includes planes.
@@
A Special Weingarten surface (SW-surface, for short) is an oriented surface $\Sigma$ immersed in $\mathbb{R}^{3}$ whose principal curvatures $\kappa_{1}\geq\kappa_{2}$ satisfy at every point of $\Sigma$ a relation of the type $\kappa_{2}=g(\kappa_{1})$, for some $g\in C^{1}(I_{g})$ with $g^{\prime}(t)<0$ everywhere, where $I_{g}=[\alpha,b)$ for some $\alpha\in\mathbb{R}$ and $b\in\mathbb{R}\cup\{+\infty\}$ and $g(\alpha)=\alpha$.
@@ 1.2 An equivalence relation among Special Weingarten surfaces
1.3 The tangency principle for Special Weingarten surfaces
@@ 2 SW-catenoids
Special Weingarten surfaces with rotational symmetry are described in details and classified in [FeMi]. The minimal-type case was previously treated in [SaTo1] and more recently in [LoPa]. Recall that, if $\Sigma$ is a surface in $\mathcal{W}_{g}$ for a given $g$ with $\alpha=0$, then $\Sigma$ is called of minimal-type. Regular surfaces invariant by rotations in a set of SW-surfaces $\mathcal{W}_{g}$ of minimal-type, when they exist, are called special catenoids or SW-catenoids.
@@
Let us consider the particular cases where the set $\mathcal{W}_{g}$ of elliptic Weingarten surfaces of minimal-type is given in terms of $g(t)=m_{0}t$, for some $m_{0}<0$, which in view of Theorem 5.3 in [FeMi] admits SW-catenoids. Let $G\in\mathcal{W}_{g}$ be a SW-catenoid that is inward oriented. After an isometry of $\mathbb{R}^{3}$ we can assume that the revolution axis is the $z$-axis, and the neck is the horizontal circle $C_{0}$ of radius $r_{0}>0$ centered at the origin. The part of $G$ above the plane $\{z=0\}$ can be parametrized as
@@ 3 The limaçon of Pascal
Denoting by $C$ and $c$ respectively the center and the radius of $\mathcal{C}$, it is easy to prove that, up to isometries of the Euclidean plane, $\mathcal{L}$ is uniquely determined by two positive parameters: $a=\mbox{dist}(A,C)$ and $c$. Moreover $\mathcal{L}$ depends continuously by $a$ and $c$ (see Fig. 4, left).
Figure 4: on the left, the construction of the limaçon of Pascal for $a=5$ and $c=2$; on the right, the limaçon of Pascal and the disks described in Lemma 3.3.
@@ 4 The graph Lemma
We denote by $D(p,r)$ the horizontal closed disk of radius $r$ and center $p$. Moreover, for any interval $I\subset\mathbb{R}$, we denote by $D(p,r)\times I$ the intersection of the solid cylinder on $D(p,r)$ with the closed slab $\{(x,y,z)\in\mathbb{R}^{3};z\in I\}$.
For convenience, let us recall a well known property about strictly convex planar curves: let $\Gamma\subset\mathbb{R}^{2}$ be a strictly convex planar curve with maximum and minimum curvature values $\Lambda$ and $\lambda$ satisfying $\Lambda>\lambda>0$. Then any circle of radius $1/\lambda$ (resp. $1/\Lambda$) tangent to $\Gamma$, with the same inward normal vector as $\Gamma$ at the tangency point, must bound a closed domain containing $\Gamma$ (resp. contained in the closed domain bounded by $\Gamma$). Moreover, the radius $\omega(\Gamma)$ of the smallest circle in $\mathbb{R}^{2}$ enclosing $\Gamma$ satisfies $1/\Lambda\leq\omega(\Gamma)\leq 1/\lambda$.
Now, let us prove a result that will be crucial in what follows.
@@
If $\Omega$ is the planar domain bounded by $\Gamma,$ then there exists $p\in\Omega$ depending on $\Gamma$ and $\Sigma$ and $r>0$ depending only on $\Gamma$ such that $\Sigma\cap(D(p,r)\times[0,+\infty))$ is a vertical graph over $D(p,r)$, with
@@
We denote by $\mathcal{O}$ the open region of $\mathbb{R}^{3}$ bounded by $\Sigma\cup\Omega$. The strategy of the proof is to find a point $q$ on $\Gamma$ such that $\Sigma\cap(\{q\}\times(0,+\infty))$ is only one point. Then, the disk $D(p,r)$ of the statement will be obtained, by reflecting points closed to $q$ with respect to vertical planes, and using the Alexandrov reflection method.
@@
In order to find $q\in\Gamma$, we apply the Alexandrov reflection method with horizontal hyperplanes $P(t):=\{z=t\}$ coming down from above. Denote by $h_{\Sigma}$ the height of $\Sigma$ above $P(0)$. The plane $P(t)$, $t>h_{\Sigma}$, does not intersect $\Sigma$. Then we let $t$ decrease. When $t<h_{\Sigma}$, we reflect the part of $\Sigma$ above $P(t)$ until there is a first contact point between $\Sigma$ and its reflection. If such a contact point does not occur for any $t>0$, then $\Sigma$ is a vertical graph over $\Omega$ and we can choose $r(\Gamma)=1/\Lambda$ and $p=q+(1/\Lambda)n(q)$, where $q$ is any point of $\Gamma$ and $n(q)$ is the inward pointing unit normal vector of $\Gamma$ at $q$.
@@
Otherwise, there will be a $20<t_{0}<h_{\Sigma}/$ 2 such that the reflected surface touches $\Sigma$ for the first time. If the intersection occurs at an interior point of $\Sigma$ we would have a contradiction with the tangency principle, hence a first touching point must belong to $\Gamma$. Let $q$ be one of such points. Then the line $\{q\}\times(0,+\infty)$ intersects $\Sigma$ exactly once, and $\{q\}\times(0,2t_{0})$ is contained in $\mathcal{O}$, as $t_{0}<h_{M}/2$. Note that the portion of $\Sigma$ above $P(t_{0})$ is a graph.
Let us now show the existence of the disk $D(p,r)$.
@@
First we notice the following fact. Let $Q$ be any vertical plane. Since $\Sigma$ is compact, we can assume that $Q\cap\Sigma=\emptyset$. Translate horizontally $Q$ towards $\Sigma$. By abuse of notation, we call again $Q$ any parallel translation of the initial plane. When $Q$ touches $\Sigma$ for the first time, keep moving $Q$ and start reflecting through $Q$ the part of $\Sigma$ left behind $Q$. In order not to have a contradiction with the tangency principle, we can continue this procedure with no contact points between $\Sigma$ and its reflection until $Q$ enters $\Gamma$ at distance at least $1/\Lambda$ from it. In fact let $p^{\prime}\in\Gamma$ behind a $Q_{v},$ where $Q_{v}$ is at distance less than $1/\Lambda$ from $p\in\Gamma$. Then the disk of radius $1/\Lambda$ tangent to $\Gamma$ at $p^{\prime}$ is contained in $\Omega$ and its center is on the side of $Q_{v}$ not containing $p$. Hence, the reflection of $p^{\prime}$ is inside the disk, then it is is inside $\Omega$.
Let $\Omega_{q}$ be the set of such $p^{\prime}$ obtained with the above construction using just the vertical planes that leave $q$ behind them. We have that $\Omega_{q}$ is a subdomain of $\Omega$ with non-empty interior and $\Omega_{q}\times(0,2t_{0})\subset\mathcal{O}$. Therefore $\Omega_{q}$ contains the desired disk. Its shape and size depend on $\Gamma$, but also on $q$. Since we are not able to predict where $q$ is, we can avoid the dependence of the point $q$ stopping the reflection earlier. More precisely we stop when $Q$ becomes tangent to $\mathcal{C}$, where $\mathcal{C}$ is defined as follows. Denote by $\mathcal{C}^{\prime}$ the horizontal circle of radius $1/\lambda$, tangent to $\Gamma$ at $q$, and enclosing $\Gamma$. Then $\mathcal{C}$ is the circle with the same center as that of $\mathcal{C}^{\prime}$ and radius equal to $1/\lambda-1/\Lambda$.
@@
The set of the reflections of $q$ through any vertical plane tangent to $\mathcal{C}$ is a limaçon of Pascal $\mathcal{L}$ in the plane $P$, with respect to the circle $\mathcal{C}$ and base point $q$. The parameters of $\mathcal{L}$ are $a=1/\lambda$ and $c=1/\lambda-1/\Lambda$, which indicates that the limaçons of Pascal obtained starting with different points $q\in\Gamma$ are congruent. As $a>c$, then $\mathcal{L}$ has two loops. By item (2) of Lemma 3.3, the smaller loop $\mathcal{L}^{\prime}$ of $\mathcal{L}$ is contained in the circle of radius $a-c=1/\Lambda$ tangent to $\Gamma$ at $q$ hence it is contained in $\Omega_{q}\cup\{q\}$.
@@
Finally, the conclusion of the statement follows from the fact that the closed domain bounded by the smaller loop of $\mathcal{L}$ contains a closed disk $D(p,r)$, for some $p\in\Omega$ and $r=r(\Gamma)$ greater or equal than $\sqrt{\frac{(a-c)^{3}}{a}}=\sqrt{\frac{\lambda}{\Lambda^{3}}}$, by item (1) of Lemma 3.3, and from similar arguments we used before, $\Sigma$ is a vertical graph over $D(p,r)$. ∎